4. Given the following information on U.S. Treasury instruments:

1-year note yield = 1.83%                       5-year note yield = 2.79%

2-year note yield = 1.99%                       6-year note yield = 3.04%

3-year note yield = 2.21%                       7-year note yield = 3.58%

4-year note yield = 2.42%                       8-year note yield = 4.18%

And non-changing premiums of 0, .11%, .22%, .39%, .52%, .64%, .75%, .88%, & .98%

a. Calculate the expected expectations yield for a (3,2,1,2) path.

b. Calculate the expected empirical yields for a (2,5,1) path.

c. Calculate the expected expectations yield for a 5-year note purchased at the beginning of year 3.

d. Calculate the expected market yield on a 4-year note purchased at the beginning of year 2.

e. Determine the expectations yield on a 5-year note purchased today.

f. Determine the yield on a 6-year Treasury note purchased today.
Answer #1

Solution: (W) Given information on Vos Treasury instruments. @ Expected expectations yield for a (3,2,1,2) Path: pure expecta= expected Emperical yield for a (2,5,1) path = (1+0.221) (1+0.0279 +0.0022) (1+0.0183+ 0.0088) - = (1.0221) (0301) (1.0271)@ The expectations yield on syear note purchated today : The expectation yield as per pure expectation theory 18: ((1+2year y